OpenStudy (moonlight93):

Microwave ovens emit microwave energy with a wavelength of 12.5 cm. What is the energy of exactly one photon of this microwave radiation? I'm on sapling and I did it couple times, but I kept getting it wrong!!! I followed all the notes I have from lecture, can someone please help thank you

OpenStudy (matt101):

Can I see your work?

OpenStudy (moonlight93):

ok, hold on

OpenStudy (moonlight93):

First I converted 12.5 cm to meters which is 0.125m , then by using wavespeed = frequency x wavelength 3x10^8= f x 0.125 f= 2400000000 then I plugged it in to this equation; E= h x f E = (6.63x10^-34)(2400000000) E= 1.5912x10^-24 I get a differ answer every time i do it, lol did I do it right this time?

OpenStudy (moonlight93):

@matt101

OpenStudy (matt101):

It's easier if you use scientific notation for these kinds of questions, that way you won't accidentally type in the wrong number of 0's into your calculator. Otherwise I agree with your process. Solving for frequency you get f = 2.4 x 10^9 Solving for energy you get E = 1.59 x 10^(-24) So yes, I agree with your answer this time!

OpenStudy (moonlight93):

Thank you. correct answer this time :)

OpenStudy (moonlight93):

@matt101

OpenStudy (moonlight93):

First I converted 12.5 cm to meters which is 0.125m , then by using wavespeed = frequency x wavelength 3x10^8= f x 0.125 f= 2400000000 then I plugged it in to this equation; E= h x f E = (6.63x10^-34)(2400000000) E= 1.5912x10^-24 I get a differ answer every time i do it, lol did I do it right this time?